Aggregate Expenditure Model:
AEo = 1,260.
c = 0.9
Determine Y.
1
Expert's answer
2016-05-13T12:36:03-0400
Aggregate Expenditure function is AE = AEo + c*Y + Io + Go + NXo.
In equilibrium AE = Y.
AEo = 1,260, c = 0.9.
As Io, Go and NXo are not provided, then:
Y = AEo + c*Y,
Y = 1,260 + 0.9Y,
0.1Y = 1,260,
Y = 12,600.